# Human 3Diffusion: Realistic Avatar Creation via Explicit 3D Consistent Diffusion Models

### v1

@@ Abstract
Creating realistic avatars from a single RGB image is an attractive yet challenging problem. Due to its ill-posed nature, recent works leverage powerful prior from 2D diffusion models pretrained on large datasets. Although 2D diffusion models demonstrate strong generalization capability, they cannot provide multi-view shape priors with guaranteed 3D consistency. We propose Human 3Diffusion: Realistic Avatar Creation via Explicit 3D Consistent Diffusion. Our key insight is that 2D multi-view diffusion and 3D reconstruction models provide complementary information for each other, and by coupling them in a tight manner, we can fully leverage the potential of both models. We introduce a novel image-conditioned generative 3D Gaussian Splats reconstruction model that leverages the priors from 2D multi-view diffusion models, and provides an explicit 3D representation, which further guides the 2D reverse sampling process to have better 3D consistency. Experiments show that our proposed framework outperforms state-of-the-art methods and enables the creation of realistic avatars from a single RGB image, achieving high-fidelity in both geometry and appearance. Extensive ablations also validate the efficacy of our design, (1) multi-view 2D priors conditioning in generative 3D reconstruction and (2) consistency refinement of sampling trajectory via the explicit 3D representation. Our code and models will be released here.
Figure 1: Given a single image of a person (top), our method Human 3Diffusion creates 3D Gaussian Splats of realistic avatars with high-fidelity geometry and texture.
@@ 1 Introduction
To address these challenges, we propose 3Diffusion: realistic avatar creation via 3D consistent Diffusion models. We design our method based on two key insights: 1) 2D multi-view diffusion models provide large-scale shape priors that can help 3D reconstruction; 2) A reconstructed 3D representation ensures 3D consistency across multi-views in 2D diffusion. Specifically, we propose a novel diffusion method, which bridges 3D Gaussian Splatting (3D-GS) [34] generation with a 2D multi-view diffusion model. At every iteration, multi-view images are denoised and reconstructed to 3D-GS to be re-rendered to continue the diffusion process. This 3D lifting during iterative sampling ensures the 3D consistency of the 2D diffusion model while leveraging a large-scale foundation model trained on billions of images. Our framework elegantly combines reconstruction methods with multi-view diffusion models. In summary, our contributions are:
@@ 5.1 Experimental Setup
Datasets. We train our model on a combined 3D human dataset [1, 3, 4, 2, 19, 25, 63, 96] compromising $\sim 6000$ high quality scans. Please refer to Sec. D.1. We evaluate qualitatively and quantitatively on Sizer [70] and more challenging IIIT [30] dataset due to extremely loose clothing of traditional custom suits. Please refer to Sec. D.2 for examples in evaluation dataset.
Implementation Details. We trained our model on 8 NVIDIA A100 GPUs over approximately 5 days. Each GPU was configured with a batch size 2 and gradient accumulations of 16 steps to achieve an effective batch size of 256. For more training details regarding hyperparameters, diffusion schedulers, etc., please refer to Sec. A.1.
Evaluation Metrics. We evaluate the geometry quality using Chamfer Distance (CD), F-score [67] (w/ threshold of $0.01m$), and Normal Consistency (NC) between the extracted mesh (Sec. A.4) and the groundtruth scan. Appearance quality is assessed by rendering the reconstructed avatar from 32 novel views with uniform azimuth and 0 elevation angle. The metrics for appearance reported include multi-scale Structure Similarity (SSIM) [76], Learned Perceptual Image Patch Similarity (LPIPS) [102], and Peak Signal to Noise Ratio (PSNR) between rendered and ground-truth views. Moreover, we report the Fréchet inception distance (FID) [20] between synthesized views and ground truth renderings, which reflects the quality and realism of the unseen regions.
@@ 5.2 Realistic Avatar from Image
Figure 3: Qualitative comparison with baselines on SIZER [70] and IIIT [30] dataset. Template-based methods (SiTH [21] and SIFU [106]) cannot reconstruct loose clothing coherently and template free methods (PIFu [54] and TripoSR [72]) tend to generate blurry texture in unseen regions due to their deterministic formulation. LGM [65] and InstantMesh [87] cannot correct 3D inconsistency from 2D multi-views hence the texture is also blurry. In contract, our method is able to reconstruct avatars with realistic textures and plausible 3D geometry in both seen and unseen region.
We compare our approach against prior methods for image-to-avatar reconstruction. including template-based [21, 54, 106] and template-free [54] human reconstruction methods, as well as general image-to-3D methods [65, 72]. To further assess performance, we also fine-tuned the state-of-the-art object reconstruction method LGM [65] and its deployed multi-view diffusion model [75] on our training data, denoted as $\text{LGM}_{\text{human}}$. Quantitative evaluations reported in Tab. 1 demonstrate that our proposed method excels in reconstructing realistic avatars with accurate geometry (CD, NC, F-score) and realistic texture (SSIM, LPIPS, PSNR, FID) from a single RGB image.
@@
TripoSR [72] $2.59$ $0.360$ $0.771$ $0.743$ $0.095$ $20.05$ $26.13$
InstantMesh [87] $2.47$ $0.338$ $0.787$ $0.900$ $0.084$ $20.54$ $26.38$
LGM [65] $3.29$ $0.275$ $0.562$ $0.892$ $0.088$ $20.11$ $24.21$
PIFu [54] $2.83$ $0.333$ $0.769$ $0.907$ $0.078$ $20.66$ $32.73$
SiTH [21] $3.92$ $0.250$ $0.735$ $0.907$ $0.074$ $20.88$ $24.76$
SIFU [106] $3.34$ $0.235$ $0.739$ $0.896$ $0.085$ $20.39$ $42.63$
$\text{LGM}_{\text{human}}$ $2.08$ $0.300$ $0.661$ $0.904$ $0.075$ $20.61$ $15.01$
Ours 1.35 0.550 0.798 0.918 0.060 21.49 9.57
Table 1: Quantitative evaluation on SIZER [70] and IIIT [30] dataset. Our method reconstructs more accurate geometry (CD, F-score, NC) and realistic textures (SSIM, LPIPS, PSNR, FID).
@@ 5.3 Ablation Studies
Figure 4: 3D reconstruction conditioned on different multi-view priors. Without our 3D-consistent sampling, the 2D diffusion model cannot generate 3D consistent multi-views (MVD, $\text{MVD}_{\text{ft}}$), leading to artifacts like floating 3D Gaussians splats. Our method obtains more consistent multi-views hence better 3D-GS and rendering.
@@
Method FID $\downarrow$
Ours w/o $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$ $12.75$
Our full model $\mathbf{9.57}$
@@
Method FID $\downarrow$
w/o $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$ $107.14$
Ours $\mathbf{83.89}$
@@
Figure 6: Visualization of reconstructed mesh and synthesized novel view of generated 3D-GS on subjects from (A) Sizer [70], (B) IIIT [30], (C) Twindom [4], (D) UBC Fashion [99], (E) GSO [16] and (F) online image "the Rock". More results are presented in Appendix C and our project page.
@@ 7 Conclusion
In this paper, we introduce Human 3Diffusion, a 3D consistent diffusion model for creating realistic avatars from single RGB images. Our key ideas are two folds: 1) Leveraging strong multi-view priors from pretrained 2D diffusion models to generate 3D Gaussian Splats, and 2) Using the reconstructed explicit 3D Gaussian Splats to refine the sampling trajectory of the 2D diffusion model which enhances 3D consistency. We carefully designed a diffusion process that synergistically combines the strengths of both 2D and 3D models. Our experiments show that our approach outperforms all previous reconstruction works in both appearance and geometry. We also extensively ablate our method which proves the effectiveness of our proposed ideas. Our code and pretrained models will be released on our Project Page to foster future research.
@@ Acknowledgements
We appreciate G.Tiwari, Y.He, Y. Xiu, Z.Liu, Z.Qiu, S.Li and others for their feedback to improve the work. This work is made possible by funding from the Carl Zeiss Foundation. This work is also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 409792180 (EmmyNoether Programme, project: Real Virtual Humans) and the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Y.Xue. G. Pons-Moll is a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. R. Marin has been supported by innovation program under the Marie Skłodowska-Curie grant agreement No 101109330.
@@ A.3 Generative 3D-GS Reconstruction Model
For sampling the 3D State $\mathbf{S}$ from $\mathbf{x}_{t}^{\text{tgt}}$, $\tilde{\mathbf{x}}_{0}^{\text{tgt}}$, $\mathbf{x}^{\text{c}}$, and $t$ (eq. 7), we adopt the time-conditioned UNet-Transformer architecture [53] due to the efficiency of convolutional layers and the scalability of transformers. For enabling the awareness of camera poses in the encoding process, we concatenate the Plücker Camera Ray Embedding $\{\mathbf{o}_{i}\times\mathbf{d}_{i},\mathbf{d}_{i}\}$ [65, 88] with the image $\mathbf{x}_{t}^{\text{tgt}}$ and $\tilde{\mathbf{x}}_{0}^{\text{tgt}}$. To enhance the control ability of context view in the 3D generation process, we additionally concatenate the clear context view ${\mathbf{x}}^{\text{c}}$ with target images $\mathbf{x}_{t}^{\text{tgt}}$ follow  [75]. This operation enables 3D dense self-attention process between the input multi-view target images and the clear context view image, provides pixel-level local conditional signal. Since the camera pose of context view $\mathbf{x}^{c}$ is unknown, we use the 00-vector as its embedding.
Figure 7: Visualization intermediate sampling steps from a Gaussian Noise ($t=1000$) to the last denoising step ($t=0$). From top to bottom: current state $\mathbf{x}^{\text{tgt}}_{t}$, estimated clear view by 2D diffusion models $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$, and corrected clear view by generated 3D Gaussian Splatting $\hat{\mathbf{x}}^{\text{tgt}}_{0}$. Our 2D diffusion model $\bm{\epsilon}_{\phi}(\circ)$ already provides strong multi-view prior at an early stage with large $t$. Our 3D reconstruction model $\mathbf{g}_{\phi}(\circ)$ can correct the inconsistency in $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$ illustrated in red circle.
@@ B.1 Qualitative Comparison
Figure 9: Qualitative comparison on IIIT [30].
@@ D.2 Evaluation Dataset
For quantitative evaluation, we use Sizer [70] and IIIT 3D human dataset [30]. In this section, we start with introducing the two used evaluation datasets, and explain why we omit the commonly used CAPE dataset [45, 50, 101] in our experiments. Finally, we provide a summary of the evaluation datasets.
@@
Why not CAPE
Unlike previous methods [21, 85, 86, 106] which evaluate on CAPE dataset [45, 50, 101], we omitted CAPE in our evaluation mainly because of the limitation in appearance variety, geometrical artifacts as well as the publicly unavailability. CAPE only contains simple clothing such as T-shirts and jeans, but no garments of loose clothing. As illustrated in Fig. 22, CAPE contains several artifacts, such as holes on the head, missing hands, and wrong mesh geometry. Moreover, CAPE doesn’t have most original scans publicly available, but only the SMPL+D fitting. Due to this reason, we cannot render the CAPE scan to RGB images at desired camera view to evaluate the appearance performance such as PSNR, SSIM, and LPIPS.
@@
We observe that the high quality datasets Sizer [70] and IIIT 3D Human [30] are unexplored for the community of 3D avatars reconstruction. In fact, Sizer [70] contains casual clothing which is suitable to evaluate performance, and IIIT [30] contains challenging texture and loose clothing which is suitable to evaluate robustness. All high quality scans in [30, 70] have no severe artifacts and are fully publicly available, which are the benefits unprovided in CAPE dataset [45, 50, 101]. By evaluating on these datasets [30, 70] and release our randomly sampled subjects which are used in our experiments, we hope the 3D avatars community can discover and benefit from them.
@@
Figure 22: Example artifacts in CAPE [45, 50, 101] dataset. Images shown here are rendered by ICON [85] due to the inaccessibility of original scans.

### v2

# Human-3Diffusion: Realistic Avatar Creation via Explicit 3D Consistent Diffusion Models

@@ Abstract
Creating realistic avatars from a single RGB image is an attractive yet challenging problem. To deal with challenging loose clothing or occlusion by interaction objects, we leverage powerful shape prior from 2D diffusion models pretrained on large datasets. Although 2D diffusion models demonstrate strong generalization capability, they cannot provide multi-view shape priors with guaranteed 3D consistency. We propose Human-3Diffusion: Realistic Avatar Creation via Explicit 3D Consistent Diffusion. Our key insight is that 2D multi-view diffusion and 3D reconstruction models provide complementary information for each other. By coupling them in a tight manner, we can fully leverage the potential of both models. We introduce a novel image-conditioned generative 3D Gaussian Splats reconstruction model that leverages the prior from 2D multi-view diffusion models, and provides an explicit 3D representation, which further guides the 2D reverse sampling process to have better 3D consistency. Experiments show that our proposed framework outperforms state-of-the-art methods and enables the creation of realistic avatars from a single RGB image, achieving high-fidelity in both geometry and appearance. Extensive ablations also validate the efficacy of our design, (1) multi-view 2D priors conditioning in generative 3D reconstruction and (2) consistency refinement of sampling trajectory via the explicit 3D representation. Our code and models will be released here.
Figure 1: Given a single image of a person (top), our method Human-3Diffusion creates 3D Gaussian Splats of realistic avatars with cloth and interacting objects with high-fidelity geometry and texture.
@@ 1 Introduction
To address these challenges, we propose Human-3Diffusion: realistic avatar creation via 3D consistent Diffusion models. We design our method based on two key insights: 1) 2D multi-view diffusion models provide large-scale shape priors that can help 3D reconstruction; 2) A reconstructed 3D representation ensures 3D consistency across multi-views in 2D diffusion. Specifically, we propose a novel diffusion method, which bridges 3D Gaussian Splatting (3D-GS) [36] generation with a 2D multi-view diffusion model. At every iteration, multi-view images are denoised and reconstructed to 3D-GS to be re-rendered to continue the diffusion process. This 3D lifting during iterative sampling ensures the 3D consistency of the 2D diffusion model while leveraging a large-scale foundation model trained on billions of images. Our framework elegantly combines reconstruction methods with multi-view diffusion models. In summary, our contributions are:
@@ 5.1 Experimental Setup
Datasets. We train our model on a combined 3D human dataset [1, 3, 4, 2, 21, 27, 65, 98] compromising $\sim 6000$ high quality scans. We evaluate qualitatively and quantitatively on CAPE [47, 52, 103], Sizer [8, 72] and IIIT [32] dataset. Please refer to Sec. D.1 and Sec. D.2 for more details.
Implementation Details. We trained our model on 8 NVIDIA A100 GPUs over approximately 5 days. Each GPU was configured with a batch size 2 and gradient accumulations of 16 steps to achieve an effective batch size of 256. For more training details regarding hyperparameters, diffusion schedulers, etc., please refer to Sec. A.1 for more details. Our model creates 3D Avatar from single images in 22.6 seconds on a A100 GPU and only consumes 11.7 GB VRAM, which allows the efficient large-scale avatar generation.
Evaluation Metrics. We evaluate the geometry quality using Chamfer Distance (CD in ${cm}$), Point-to-Surface Distance (P2S in ${cm}$), F-score [69] (w/ threshold of $0.01m$), and Normal Consistency (NC) between the extracted mesh (Sec. A.4) and the groundtruth scan. Appearance quality is assessed by rendering the reconstructed avatar from 32 novel views with uniform azimuth and 0 elevation angle. The metrics for appearance reported include multi-scale Structure Similarity (SSIM) [78], Learned Perceptual Image Patch Similarity (LPIPS) [104], and Peak Signal to Noise Ratio (PSNR) between rendered and ground-truth views. Moreover, we report the Fréchet inception distance (FID) [22] which reflects the quality and realism of the unseen regions.
@@ 5.2 Realistic Avatar from Image
Figure 3: Qualitative comparison with baselines. Recent avatar reconstruction works ICON [87], ECON [88], SiTH [23] and SIFU [108]) cannot reconstruct loose clothing coherently. Additionally, SiTH and SIFU generate blurry texture in unseen regions due to their deterministic formulation of regressing 3D avatar directly from single RGB imagse. In contract, our method is able to reconstruct avatars with realistic textures and plausible 3D geometry in both seen and unseen region.
We compare our approach against prior methods for image-to-avatar reconstruction, including template-based [18, 23, 87, 88, 56, 108], template-free [56] human reconstruction methods, as well as general image-to-3D methods [67, 74, 89]. To further assess performance, we also fine-tuned the state-of-the-art object reconstruction method LGM [67] and its deployed multi-view diffusion model [77] on our training data, denoted as $\text{LGM}_{\text{human}}$. Quantitative evaluations reported in Tab. 1 demonstrate that our proposed method excels in reconstructing realistic avatars with accurate geometry (CD, NC, F-score) and realistic texture (SSIM, LPIPS, PSNR, FID) from a single RGB image.
@@
PIFu [56] $2.75$ $2.68$ $0.359$ $0.778$ $0.909$ $0.077$ $21.06$ $29.57$
SiTH [23] $4.00$ $4.00$ $0.257$ $0.749$ $0.907$ $0.073$ $20.00$ $22.33$
SIFU [108] $3.50$ $3.50$ $0.273$ $0.760$ $0.0.900$ $0.081$ $20.73$ $40.75$
$\text{LGM}_{\text{human}}$ $3.44$ $3.44$ $0.272$ $0.560$ $0.893$ $0.088$ $20.56$ $14.22$
FoF [18] $5.43$ $5.34$ $0.183$ $0.683$
ICON [87] $3.88$ $3.94$ $0.244$ $0.749$
ECON [88] $2.83$ $3.61$ $0.291$ $0.767$
Ours 1.41 1.37 0.557 0.791 0.916 0.058 21.61 8.45
Table 1: Quantitative evaluation on CAPE [47], SIZER [72], and IIIT [32] dataset. Our method can perform better reconstruction in terms of more accurate geometry (CD, P2S, F-score, NC) and realistic textures (SSIM, LPIPS, PSNR, FID).
@@ 5.3 Ablation Studies
Figure 4: 3D reconstruction conditioned on different multi-view priors. Without our 3D-consistent sampling, the 2D diffusion model cannot generate 3D consistent multi-views (MVD, $\text{MVD}_{\text{ft}}$), leading to artifacts like floating 3D Gaussians splats.
@@
Method PSNR $\uparrow$
Ours w/o $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$ $20.98$
Our full model $\mathbf{21.49}$
@@
Method PSNR $\uparrow$
w/o $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$ $14.45$
Ours $\mathbf{16.12}$
@@
Figure 6: Visualization of reconstructed mesh and synthesized novel view of generated 3D-GS on subjects from Sizer [72], RenderPeople [2], Twindom [4], UBC Fashion [101], GSO [17] and online image. More results are presented in Appendix C and our project page.
@@ 7 Conclusion
In this paper, we introduce Human-3Diffusion, a 3D consistent diffusion model for creating realistic avatars from single RGB images. Our key ideas are two folds: 1) Leveraging strong multi-view priors from pretrained 2D diffusion models to generate 3D Gaussian Splats, and 2) Using the reconstructed explicit 3D Gaussian Splats to refine the sampling trajectory of the 2D diffusion model which enhances 3D consistency. We carefully designed a diffusion process that synergistically combines the strengths of both 2D and 3D models. Our experiments show that our approach outperforms all previous reconstruction works in both appearance and geometry. We also extensively ablate our method which proves the effectiveness of our proposed ideas. Our code and pretrained models will be released on our Project Page to foster future research.
@@ Acknowledgements
We appreciate G.Tiwari, Y.He, Y. Xiu, Z.Liu, Z.Qiu, S.Li and others for their feedback to improve the work. This work is made possible by funding from the Carl Zeiss Foundation. This work is also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 409792180 (EmmyNoether Programme, project: Real Virtual Humans) and the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Y.Xue. R. Marin has been supported by innovation program under the Marie Skłodowska-Curie grant agreement No 101109330. G. Pons-Moll is a member of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645.
@@ A.3 Generative 3D-GS Reconstruction Model
For sampling the 3D State $\mathbf{S}$ from $\mathbf{x}_{t}^{\text{tgt}}$, $\tilde{\mathbf{x}}_{0}^{\text{tgt}}$, $\mathbf{x}^{\text{c}}$, and $t$ (eq. 7), we adopt the time-conditioned UNet-Transformer architecture [55] due to the efficiency of convolutional layers and the scalability of transformers. For enabling the awareness of camera poses in the encoding process, we concatenate the Plücker Camera Ray Embedding $\{\mathbf{o}_{i}\times\mathbf{d}_{i},\mathbf{d}_{i}\}$ [67, 90] with the image $\mathbf{x}_{t}^{\text{tgt}}$ and $\tilde{\mathbf{x}}_{0}^{\text{tgt}}$. To enhance the control ability of context view in the 3D generation process, we additionally concatenate the clear context view ${\mathbf{x}}^{\text{c}}$ with target images $\mathbf{x}_{t}^{\text{tgt}}$ following [77]. This operation enables 3D dense self-attention process between the input multi-view target images and the clear context view image, provides pixel-level local conditional signal. Since the camera pose of context view $\mathbf{x}^{c}$ is unknown, we use the 00-vector as its embedding.
Figure 7: Visualization of intermediate sampling steps from a Gaussian Noise ($t=1000$) to the last denoising step ($t=0$). From top to bottom: current state $\mathbf{x}^{\text{tgt}}_{t}$, estimated clear view by 2D diffusion models $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$, and corrected clear view by generated 3D Gaussian Splatting $\hat{\mathbf{x}}^{\text{tgt}}_{0}$. Our 2D diffusion model $\bm{\epsilon}_{\phi}(\circ)$ already provides strong multi-view prior at an early stage with large $t$. Our 3D reconstruction model $\mathbf{g}_{\phi}(\circ)$ can correct the inconsistency in $\tilde{\mathbf{x}}^{\text{tgt}}_{0}$ illustrated in red circle.
@@ B.1 Qualitative Comparison
Figure 9: Qualitative comparison on Sizer [72] and IIIT [32].
@@ D.2 Evaluation Dataset
For quantitative evaluation, we use Sizer [72] and IIIT 3D human dataset [32]. In this section, we start with introducing the two used evaluation datasets, and explain why we omit the commonly used CAPE dataset [47] in our experiments. Finally, we provide a summary of the evaluation datasets.
CAPE
Cape [47] is a popular evaluation dataset which is widely used by previous methods [23, 87, 88, 108]. We observe that CAPE has limitation in appearance variety, geometrical artifacts as well as the publicly unavailability. CAPE only contains simple clothing such as T-shirts and jeans, but no garments of loose clothing. As illustrated in Fig. 22, CAPE contains several artifacts, such as holes on the head, missing hands, and wrong mesh geometry. We manually removed the noise and artifacts in scan to serve as our evaluation dataset. Moreover, CAPE doesn’t have most original scans publicly available, but only the SMPL+D fitting. Due to this reason, we cannot render the CAPE scan to RGB images at desired camera view to evaluate the appearance performance such as PSNR, SSIM, and LPIPS. For above mentioned reasons, we also propose to use Sizer [72] and IIIT [32] which are publicly available.
@@
We observe that the high quality datasets Sizer [72] and IIIT 3D Human [32] are unexplored for the community of 3D avatars reconstruction. In fact, Sizer [72] contains casual clothing which is suitable to evaluate performance, and IIIT [32] contains challenging texture and loose clothing which is suitable to evaluate robustness. All high quality scans in [32, 72] have no severe artifacts and are fully publicly available, which are the benefits unprovided in CAPE dataset [47]. By evaluating on these datasets [32, 72] and release our randomly sampled subjects which are used in our experiments, we hope the 3D avatars community can discover and benefit from them.
@@
Figure 22: Example artifacts in CAPE [47] dataset. Images shown here are rendered by ICON [87] due to the inaccessibility of original scans.